OpenStudy (albert0898):

If f(x) = radical(x) and f[g(x)] = 2 radical(x) , then g(x) = Answer is 4x, but how...?

OpenStudy (mathmale):

First, move forward to verify your answer, g(x)=4x. Write f(x)=radical(x). Replace the x in f(x) with g(x) to obtain f(g(x)). Replace the x in radical(x) with 4x. Simplify the result. If you obtain 2radical(x), your choice of g(x) is correct (that is, 4x is correct).

OpenStudy (albert0898):

How would I be able to figure it out without knowing the answer?

OpenStudy (mathmale):

How would we move backwards? if f(g(x))=2 radical (x), re-write that '2' as radical (4). Then f(g(x)) = 2 radical (x) = radical (4)*radical (x)=?

OpenStudy (mathmale):

As y ou'll see above, I planned to and have taken you in both directions.

OpenStudy (mathmale):

Can you combine radical(4) and radical(x) into just one radical?

OpenStudy (albert0898):

So f(g(x)) = radical(4x)

OpenStudy (mathmale):

\[\sqrt{4}*\sqrt{x}=\sqrt{?}\]

OpenStudy (albert0898):

\[\sqrt{4x}\]

OpenStudy (zarkon):

\[f(x) = \sqrt{x}\] \[f(g(x))=2\sqrt{x}\] thus \[\sqrt{g(x)}=2\sqrt{x}\] square both sides

OpenStudy (albert0898):

Why did you put a radical around the left portion? @Zarkon

OpenStudy (mathmale):

Albert: referring to your \[\sqrt{4x}\]... simplify this and then compare your result to the original problem statement. Do you see similarities or differences there?

OpenStudy (albert0898):

I get \[2\sqrt{x}\] That's the similarity that I see

OpenStudy (mathmale):

Note that "radical(x)" in this context means sqrt(x). Yes. Have you compared that to the original problem statement?

OpenStudy (mathmale):

f[g(x)] = 2 radical(x)

OpenStudy (albert0898):

Yes, so now I have to square both sides. But how shall I square the left portion?

OpenStudy (mathmale):

Why do you have to square both sides? Zarkon has suggested that you do this, yes, but can you yourself see the reason for doing so?

OpenStudy (albert0898):

Other than getting rid of the radical, then no, I do not see a reason for doing so

OpenStudy (mathmale):

Getting rid of the radical is the whole idea. Then you're left with g(x) on the left side and 4x on the right side.

OpenStudy (mathmale):

...this goes to show that g(x) is indeed = to 4x. g(x)=4x.

OpenStudy (albert0898):

But how do we get rid of the f in f(g(x)). That's what I'm confused about now

OpenStudy (mathmale):

I was confused at first also. Zarkon correctly pointed out that f(x) = sqrt(x). So, he started with f(g(x)) and replaced the f( ) with sqrt( ).

OpenStudy (mathmale):

so he / we ended up with Sqrt(g(x)) = 2Sqrt(x). Square both sides of that. What do you obtain?

OpenStudy (albert0898):

OH! Wow, this question has a bunch of twists and turns. Wow, I didn't even think of that. Now it all makes sense!

OpenStudy (mathmale):

I'm so glad. Let's thank Zarkon.

OpenStudy (albert0898):

We obtain 4x because in order to get rid of the radicals we have to raise both sides to the 2nd power.

OpenStudy (mathmale):

Exactly.

OpenStudy (albert0898):

Thank you @Zarkon! Thank you @mathmale!

OpenStudy (mathmale):

:) :)